# Training on test proteins improves fitness, structure, and function prediction

### v1

@@ Appendix A Experimental details
In this section, we describe the experimental details for the three downstream tasks considered in this work: protein fitness prediction (Section A.1), protein structure prediction (Section A.2), and protein function prediction (Section A.3). Each subsection describes the application of test-time training (TTT) to the respective models, along with details on the datasets, metrics, and models. Table 4 additionally summarizes the hyperparameters used for the application of TTT to individual models.
A.1 Protein fitness prediction
A.1.1 Datasets
ProteinGym.
ProteinGym is the standard benchmark for protein fitness prediction (Notin et al., 2024). The latest, second version of the dataset includes 217 deep mutation scanning experiments (DMSs) across different proteins. In total, the dataset contains 2.5 mutants with annotated ground-truth fitness. Since ProteinGym does not contain a data split for the zero-shot setup, employed in this work, we use the whole dataset as the test set.
MaveDB dataset.
Figure 6: Comparison of the standard ProteinGym dataset with the MaveDB dataset constructed in this work. A) MaveDB, mined from Esposito et al. (2019), includes novel assays even after filtering to ensure distinct proteins from the comprehensive ProteinGym dataset. This is largely because most MaveDB assays post-filtering date to 2024, whereas the latest assays in ProteinGym date to 2023. B, C, D) MaveDB is of sufficient quality for model evaluation. Representative baselines, ESM2 and SaProt with both 35 million and 650 million parameters, evaluated on ProteinGym generalize effectively to MaveDB, following a similar distribution of predictions. Panel D illustrates the random subset of 50 proteins used for hyperparameter tuning for fitness prediction. Each point in the plots represents one protein and shows the Spearman correlation averaged across all assays corresponding to the protein (typically one assay per protein). The box plots standardly depict quartiles, medians, and outliers.
The described methodology resulted in the MaveDB dataset comprising 676 assays (out of 1178 in the entire MaveDB) with experimental fitness annotations. This corresponds to 483 unique protein sequences and 867 thousand mutations in total. The large size of the dataset, despite the comprehensiveness of ProteinGym containing 217 assays, can be attributed to the fact that many assays in MaveDB were released after the ProteinGym construction (Figure 6A). To ensure the quality of the constructed MaveDB dataset, we validated that representative baselines from ProteinGym generalize to the new assays, following a similar distribution of predictions (Figure 6B,C). Finally, for efficiently tuning hyper-parameters for fitness prediction models we sampled 50 random proteins (Figure 6D), corresponding to 83 assays and collectively 134 thousand variants.
A.1.2 Metrics
Spearman by phenotype.
When computing Spearman correlations, we follow the evaluation protocol proposed in ProteinGym (Notin et al., 2024). First, for each protein, we compute Spearman correlation scores between the predicted ranks of mutations and their corresponding labels. Then, we average the scores across five categories of assayed phenotypes, measuring the effects of introduced mutations: protein catalytic activity (“Activity”), binding affinity to a target (“Binding”), protein expression levels in a cell (“Expression”), organism growth rate (“Organismal Fitness”), and protein thermostability (“Stability”).
Avg. Spearman.
We refer to the mean score across the five phenotype categories as “Avg. Spearman”. We report the “Avg. Spearman” metric as the mean and standard deviation across five random seeds (Section 4.1, Table 5).
Spearman by MSA Depth.
Following (Notin et al., 2024), we split the performance by the depth of available multiple sequence alignment (MSA), i.e., the number of homologous sequences available, as provided in ProteinGym: “Low depth”, “Medium depth”, and “High depth”, and report the Spearman correlation for each subset individually (Table 6). Specifically, the MSA depth categories in ProteinGym are determined using the following thresholds from Hopf et al. (2017): “Low” is defined as $N_{eff}/L<1$, “Medium” as $1001<N_{eff}/L<$ 100, and “High” as $N_{eff}/L>100$, where $N_{eff}$ represents the normalized number of effective sequences in the MSA, and $L$ is the sequence length covered in the MSA.
A.1.3 Models
ESM2.
The ESM2 model is a bidirectional, BERT-like (Devlin, 2018) transformer trained on millions of protein sequences using masked modeling (Lin et al., 2023). The goal of protein fitness prediction is to predict the effects of mutations, and protein language models are often adapted to this task using zero-shot transfer via log odds ratio (Notin et al., 2024; Meier et al., 2021). Specifically, for a given single- or multi-point mutation, where certain amino acids $T$ are substituted from $x_{i}$ to $x^{m}_{i}$ for each $i\in T$, the fitness prediction via the log odds ratio is defined as:
$\sum_{i\in T}{\log{p(x^{m}_{i}|x_{\setminus i})}-\log{p(x_{i}|x_{\setminus i})}},$ (3)
where the sum iterates over mutated positions $i\in T$ with $p(x^{m}_{i}|x_{\setminus i})$ and $p(x_{i}|x_{\setminus i})$ denoting the predicted probabilities of the mutated amino acid and the original one (i.e., wild type), respectively. The conditionals $x_{\setminus i}$ indicate that the input sequence to the model has the position $i$ masked. In this setup, the native (unmutated) sequence, where $T=\emptyset$, has a predicted fitness of 0. Mutations with negative values represent favorable mutations, while positive values correspond to disruptive mutations. We follow the ProteinGym benchmark [^4] https://github.com/OATML-Markslab/ProteinGym and use this formula (Notin et al., 2024) to evaluate the fitness prediction capabilities of ESM2. We use the implementation of ESM2 from ProteinGym.
ESM2 + TTT.
ESM2 can be straightforwardly enhanced with test-time training. Specifically, we treat the transformer encoder as the backbone $f$, and the language modeling head, which projects token embeddings to amino acid probabilities, as the pre-training head $g$. The log odds ratio given by Equation 3 serves as the task-specific head $h$, which in this case involves the pre-training head $g$ that predicts log probabilities. Overall, we apply TTT to the pre-trained ESM2 model and, after a pre-defined number of self-supervised fine-tuning steps, score mutations using Equation 3. During TTT we fine-tune all parameters in $g\circ f$ end-to-end except for token and position embeddings.
SaProt.
We also experiment with the state-of-the-art fitness prediction model, SaProt (Su et al., 2023). SaProt builds off the ESM2 model but incorporates structural information from predicted protein structures. Specifically, SaProt uses the same transformer architecture but expands its vocabulary by combining the 20 standard amino acid tokens with 20 structural tokens from the 3Di vocabulary, increasing the total alphabet size to 400. The 3Di tokens capture the geometry of the protein backbone and are generated using VQ-VAE (Razavi et al., 2019), which projects continuous geometric information into discrete tokens and was trained as part of the Foldseek method (van Kempen et al., 2022).
Since SaProt is also a protein language model, it also uses Equation 3 to score variants. However, please note that SaProt, as implemented in ProteinGym (Notin et al., 2024), uses a slightly different version of the log odds ratio. In SaProt, the conditions in the log probabilities in Equation 3 are replaced with $x_{\setminus T}$ instead of $x_{\setminus i}$, not assuming the independence of substitutions. During TTT, we only mask sequential information and leave the structural part of the tokens unchanged, reflecting the original pre-training setup. We use the implementation of SaProt from ProteinGym4.
SaProt + TTT.
Since the architecture of SaProt is based on ESM2, the TTT components $f$, $g$, and $h$ remain the same. It means that test-time training can be applied to the model in the same way as in the case of ESM2 + TTT discussed above.
A.2 Protein structure prediction
A.2.1 Datasets
@@
To evaluate the capabilities of TTT on protein folding, we employ the CAMEO validation and test sets as described in Lin et al. (2023). Specifically, the validation set was obtained by querying the CAMEO (Continuous Automated Model Evaluation) web server [^5] https://www.cameo3d.org/modeling (Robin et al., 2021) for entries between August 2021 and January 2022, while the CAMEO test set consists of entries from April 1, 2022, to June 25, 2022. Most of the entries in the CAMEO sets are predicted with high accuracy and confidence (Lin et al., 2023). Therefore, we subselected the challenging validation and test sets where TTT is relevant.
Specifically, we applied two criteria: (1) preserving entries with ESMFold pLDDT scores below 70 to filter out high-confidence predictions (Jumper et al., 2021), and (2) selecting entries with ESM2 perplexity scores greater than or equal to 6, ensuring that the predictions are challenging due to poor sequence understanding rather than other factors. Additionally, most structures with perplexity scores below 6 are already associated with high-confidence predictions (Figure S5 in Lin et al. (2023)). After filtering, the resulting challenging validation and test sets consist of 27 (out of 378) and 18 (out of 194) targets, respectively. The vast majority of the remaining structures have accurate ESMFold structure predictions.
A.2.2 Metrics
@@ A.2.3 Models
ESMFold + TTT.
Since ESM2 backbone of ESMFold was pre-trained in a self-supervised masked modeling regime, the application of TTT to ESMFold is straightforward. We treat ESM2 as the backbone $f$, the language modeling head predicting amino acid classes from their embeddings as the self-supervised head $g$, and the folding trunk along with the structure modules as the downstream task head $h$. After each TTT step, we run $h\circ f$ to compute the pLDDT scores, which allows us to estimate the optimal number of TTT steps for each protein based on the highest pLDDT score.
Since the backbone $f$ is given by the ESM2 model containing 3 billion parameters, we apply LoRA (Hu et al., 2021) to all matrices involved in self-attention. This enables fine-tuning ESMFold + TTT on a single GPU.
@@
Since ESMFold is a regression model, it only predicts one solution and does not have a straightforward mechanism of sampling multiple structure predictions. Nevertheless, the authors of ESMFold propose a way to sample multiple candidates (Section A.3.2 in Lin et al. (2023)). To sample more solutions, the masking prediction (ME) method randomly masks 15% (same ratio as during masked language modeling pre-training) of the amino acids embeddings before passing them to the structure prediction block. Selecting the solution with the highest pLDDT may lead to improved predicted structure. Since sampling multiple solutions with ESMFold + ME and selecting the best one via pLDDT is analogous to ESMFold + TTT, we employ the former as a baseline, running the method for the same number of step.
@@
ESM3 + TTT.
[…]
[] Refer to caption
Figure 7: Average running time of ESM3 + TTT (50 steps) on the CAMEO validation set using a single NVIDIA A100 GPU.
We treat the transformer encoder of ESM3 as $f$, the language modeling head decoding amino acid classes as $g$, and the VQ-VAE decoder, which maps structure tokens to the 3D protein structure, as $h$. During the TTT steps, we train the model to unmask a protein sequence while keeping the structural track fully padded. During the inference, we provide the model with a protein sequence and run it to unmask the structural tokens, which are subsequently decoded with the VQ-VAE decoder. After each TTT step, we run $h\circ f$ to compute the pLDDT scores, which allows us to estimate the optimal number of TTT steps for each protein based on the highest pLDDT score. We choose the optimal hyperparameters by maximizing the difference in TM-score after and before applying TTT across the validation dataset.
Despite the fact that the model contains 1.4 billion parameters, even without using LoRA, ESM3 + TTT can be fine-tuned on a single NVIDIA A100 GPU. Therefore, we do not employ LoRA for fine-tuning ESM3. Figure 7 shows that ESM3 + TTT can be loaded in VRAM and run for 50 steps in just 10 seconds, with an additional 10 seconds overhead for pLDDT estimation. For comparison, the base ESM3 model can be run on the same hardware in 1 second.
@@
A.3 Protein function prediction
A.3.1 Datasets
@@ A.3.2 Metrics
The TPS substrate prediction problem is a 12-class multi-label classification task over possible TPS substrates. Therefore, we assess the quality of the predictions using standard multi-label classification metrics such as mean average precision (mAP) and area under the receiver operating characteristic curve (AUROC) averaged across individual classes. These metrics were used in the original TerpeneMiner paper (Samusevich et al., 2024). We report the performance by averaging the metric values concatenated across all validation folds from the 5-fold cross-validation schema.
@@
A.3.3 Models
TerpeneMiner.
TerpeneMiner is a state-of-the-art method for the classification of terpene synthase (TPS) substrates (Samusevich et al., 2024). The model consists of two parallel tracks. Given a protein sequence, TerpeneMiner first computes its ESM-1v embedding (Meier et al., 2021) and a vector of similarities to the functional domains of proteins from the training dataset, based on unsupervised domain segmentation of AlphaFold2-predicted structures (Jumper et al., 2021). The ESM-1v embedding and the similarity vector are then concatenated and processed by a separately trained random forest, which predicts TPS substrate class probabilities.
In our experiments, we use the “PLM only” version of the model, which leverages only ESM-1v embeddings (PLM stands for protein language model). This version exhibits a minor performance decrease compared to the full model but exactly follows a Y-shaped architecture, allowing us to validate the effectiveness of test-time training for predicting TPS substrates. We use the implementation of TerpeneMiner available at the official GitHub page  [^8] https://github.com/pluskal-lab/TerpeneMiner.
TerpeneMiner + TTT.
When applying TTT to TerpeneMiner, we treat the frozen ESM-1v model as a backbone $f$, its language modeling head as a self-supervised head $g$, and the random forest classifying TPS substrates as a downstream supervised head $h$.

### v2

@@ Appendix E Experimental details
In this section, we describe the proposed benchmark suite for the three customization tasks considered in this work: protein structure prediction (Section˜E.1), protein fitness prediction (Section˜E.2), and protein function prediction (Section˜E.3). Each subsection describes the application of ProteinTTT to the respective models, along with details on the data, metrics, and models. Table˜A3 additionally summarizes the hyperparameters used for the application of ProteinTTT to individual models.
E.1 Protein structure prediction
E.1.1 Datasets
@@
To evaluate the capabilities of ProteinTTT on protein structure prediction, we employ the CAMEO validation and test sets as described in Lin et al. (2023). Specifically, the validation set was obtained by querying the CAMEO (Continuous Automated Model Evaluation) web server [^4] https://www.cameo3d.org/modeling (Robin et al., 2021) for entries between August 2021 and January 2022, while the CAMEO test set consists of entries from April 1, 2022, to June 25, 2022. Most of the entries in the CAMEO sets are predicted with high accuracy and confidence (Lin et al., 2023). Therefore, we subselect the challenging validation and test sets where customization with ProteinTTT is suitable.
Specifically, we apply two standard criteria: (1) preserving entries with ESMFold pLDDT scores below 70 to filter out high-confidence predictions (Jumper et al., 2021), and (2) selecting entries with ESM2 perplexity scores greater than or equal to 6, ensuring that the predictions are challenging due to poor sequence understanding rather than other factors. Additionally, most structures with perplexity scores below 6 are already associated with high-confidence predictions (Figure S5 in Lin et al. (2023)). After filtering, the resulting challenging validation and test sets consist of 27 (out of 378) and 18 (out of 194) targets, respectively.
E.1.2 Metrics
@@ E.1.3 Models
ESMFold + ProteinTTT.
Since the ESM2 backbone of ESMFold was pre-trained in a self-supervised masked modeling regime, the application of ProteinTTT to ESMFold is straightforward. We treat ESM2 as the backbone $f$, the language modeling head predicting amino acid classes from their embeddings as the self-supervised head $g$, and the folding trunk along with the structure modules as the downstream task head $h$. After each ProteinTTT step, we run $h\circ f$ to compute the pLDDT scores, which allows us to estimate the optimal number of customization steps for each protein based on the highest pLDDT score.
Since the backbone $f$ is given by the ESM2 model containing 3 billion parameters, we apply LoRA (Hu et al., 2021) to all matrices involved in self-attention. This enables fine-tuning ESMFold + ProteinTTT on a single GPU.
@@
Since ESMFold is a regression model, it only predicts one solution and does not have a straightforward mechanism for sampling multiple structure predictions. Nevertheless, the authors of ESMFold propose a way to sample multiple candidates (Section A.3.2 in Lin et al. (2023)). To sample more predictions, the masking prediction (ME) method randomly masks 15% (same ratio as during masked language modeling pre-training) of the amino acids before passing them to the language model. Selecting the solution with the highest pLDDT may lead to improved predicted structure. Since sampling multiple solutions with ESMFold + ME and selecting the best one via pLDDT is analogous to ESMFold + ProteinTTT, we employ the former as a baseline, running the method for the same number of steps.
@@
ESM3 + ProteinTTT.
We treat the Transformer encoder of ESM3 as $f$, the language modeling head decoding amino acid classes as $g$, and the VQ-VAE decoder, which maps structure tokens to the 3D protein structure, as $h$. During the customization steps, we train the model to unmask a protein sequence while keeping the structural track fully padded. During the inference, we provide the model with a protein sequence and run it to unmask the structural tokens, which are subsequently decoded with the VQ-VAE decoder. After each customization step, we run $h\circ f$ to compute the pLDDT scores, which allows us to estimate the optimal number of customization steps for each protein based on the highest pLDDT score. We choose the optimal hyperparameters by maximizing the difference in TM-score after and before applying ProteinTTT across the validation dataset.
Despite the fact that the model contains 1.4 billion parameters, even without using LoRA, ESM3 + ProteinTTT can be fine-tuned on a single NVIDIA A100 GPU. Therefore, we do not employ LoRA for fine-tuning ESM3, while this can also be possible.
@@
HelixFold-Single.
HelixFold-Single is an MSA-free protein structure prediction model that combines representations from a pretrained protein language model with adapted AlphaFold2 geometric modules (EvoformerS and Structure) to directly predict atomic coordinates (Fang et al., 2023). We use the official implementation [^7] https://github.com/PaddlePaddle/PaddleHelix/tree/dev/apps/protein_folding/helixfold-single
HelixFold-Single + ProteinTTT.
HelixFold-Single shares the main concept with ESMFold, and we combine it with ProteinTTT in the same way as in ESMFold + ProteinTTT.
E.2 Protein fitness prediction
E.2.1 Datasets
ProteinGym.
ProteinGym [^8] https://github.com/OATML-Markslab/ProteinGym is the standard benchmark for protein fitness prediction (Notin et al., 2024). The latest, second version of the dataset includes 217 deep mutation scanning experiments (DMSs) across different proteins. We focus on the well-established zero-shot setup of the benchmark and do not experiment with the supervised setup, as it has not yet been fully incorporated into the official codebase at the time of this study. In total, the dataset contains 2.5M mutants with annotated ground-truth fitness. Since ProteinGym does not contain a data split for the zero-shot setup, employed in this work, we use the whole dataset as the test set.
MaveDB dataset.
Figure A3: Comparison of the standard ProteinGym dataset with the MaveDB dataset constructed in this work. A) MaveDB, mined from Esposito et al. (2019), includes novel assays even after filtering to ensure distinct proteins from the comprehensive ProteinGym dataset. This is largely because most MaveDB assays post-filtering date to 2024, whereas the latest assays in ProteinGym date to 2023. B, C, D) MaveDB is of sufficient quality for model evaluation. Representative baselines, ESM2 and SaProt with both 35 million and 650 million parameters, evaluated on ProteinGym generalize effectively to MaveDB, following a similar distribution of predictions. Panel D illustrates the random subset of 50 proteins used for hyperparameter tuning for fitness prediction. Each point in the plots represents one protein and shows the Spearman correlation averaged across all assays corresponding to the protein (typically one assay per protein). The box plots standardly depict quartiles, medians, and outliers.
The described methodology resulted in the MaveDB dataset comprising 676 assays (out of 1178 in the entire MaveDB) with experimental fitness annotations. This corresponds to 483 unique protein sequences and 867 thousand mutations in total. The large size of the dataset, despite the comprehensiveness of ProteinGym containing 217 assays, can be attributed to the fact that many assays in MaveDB were released after the ProteinGym construction (Figure˜A3A). To ensure the quality of the constructed MaveDB dataset, we validated that representative baselines from ProteinGym generalize to the new assays, following similar distributions of predictions (Figure˜A3B,C). Finally, for efficiently tuning hyperparameters for fitness prediction models, we sampled 50 proteins (Figure˜A3D), corresponding to 83 assays comprising 134 thousand variants.
E.2.2 Metrics
Spearman by phenotype.
When computing Spearman correlations, we follow the evaluation protocol proposed in ProteinGym (Notin et al., 2024). First, for each protein, we compute Spearman correlation scores between the predicted ranks of mutations and their corresponding labels. Then, we average the scores across five categories of assayed phenotypes, measuring the effects of mutations: catalytic activity (“Activity”), binding affinity to a target (“Binding”), protein expression levels in a cell (“Expression”), organism growth rate (“Organismal Fitness”), and protein thermostability (“Stability”).
Avg. Spearman.
We refer to the mean score across the five phenotype categories as “Avg. Spearman”. We report the “Avg. Spearman” metric as the mean and standard deviation across five random seeds (Table˜4, Table˜A4).
Spearman by MSA Depth.
Following (Notin et al., 2024), we split the performance by the depth of available multiple sequence alignment (MSA), i.e., the number of homologous sequences available, as provided in ProteinGym: “Low depth”, “Medium depth”, and “High depth”, and report the Spearman correlation for each subset individually (Table˜A4). Specifically, the MSA depth categories in ProteinGym are determined using the following thresholds from Hopf et al. (2017): “Low” is defined as $N_{eff}/L<1$, “Medium” as $1<N_{eff}/L<100$, and “High” as $N_{eff}/L>100$, where $N_{eff}$ represents the normalized number of effective sequences in the MSA, and $L$ is the sequence length covered in the MSA.
E.2.3 Models
ESM2.
The ESM2 model is a bidirectional, BERT-like (Devlin, 2018) Transformer trained on millions of protein sequences using masked modeling (Lin et al., 2023). The goal of protein fitness prediction is to predict the effects of mutations, and PLMs are often adapted to this task using zero-shot transfer via log odds ratio (Notin et al., 2024; Meier et al., 2021). Specifically, for a given single- or multi-point mutation, where certain amino acids $T$ are substituted from $x_{i}$ to $x^{m}_{i}$ for each $i\in T$, the fitness prediction via the log odds ratio is defined as:
$\sum_{i\in T}{\Bigl(\log{p(x^{m}_{i}|x_{\setminus i})}-\log{p(x_{i}|x_{\setminus i})}\Bigr)},$ (5)
where the sum iterates over mutated positions $i\in T$ with $p(x^{m}_{i}|x_{\setminus i})$ and $p(x_{i}|x_{\setminus i})$ denoting the predicted probabilities of the mutated amino acid and the original one (i.e., wild type), respectively. The conditionals $x_{\setminus i}$ indicate that the input sequence to the model has the position $i$ masked. In this setup, the native (unmutated) sequence, where $T=\emptyset$, has a predicted fitness of 0. Mutations with negative values represent favorable mutations, while positive values correspond to disruptive mutations. We follow the ProteinGym benchmark and use this formula (Notin et al., 2024) to evaluate the fitness prediction capabilities of ESM2. We use the implementation of ESM2 from ProteinGym.
ESM2 + ProteinTTT.
[…]
SaProt.
We also experiment with a structure-aware protein language model, SaProt (Su et al., 2023). SaProt builds off the ESM2 model but incorporates structural information from predicted protein structures. Specifically, SaProt uses the same Transformer architecture but expands its vocabulary by combining the 20 standard amino acid tokens with 20 structural tokens from the 3Di vocabulary, increasing the total alphabet size to 400. The 3Di tokens capture the geometry of the protein backbone and are generated using VQ-VAE (Razavi et al., 2019), which projects continuous geometric information into discrete tokens and was trained as part of the Foldseek method (van Kempen et al., 2022).
Since SaProt is also a protein language model, it also uses Equation˜5 to score variants. However, please note that SaProt, as implemented in ProteinGym (Notin et al., 2024), uses a slightly different version of the log odds ratio. In SaProt, the conditions in the log probabilities in Equation˜5 are replaced with $x_{\setminus T}$ instead of $x_{\setminus i}$, not assuming the independence of substitutions. During customization with ProteinTTT, we only mask sequential information and leave the structural part of the tokens unchanged, reflecting the original pre-training setup. We use the implementation of SaProt from ProteinGym8.
SaProt + ProteinTTT.
Since the architecture of SaProt is based on ESM2, the ProteinTTT components $f$, $g$, and $h$ remain the same. It means that customization can be applied to the model in the same way as in the case of ESM2 + ProteinTTT discussed above.
ProSST.
Finally, we experiment with the state-of-the-art fitness predictor, ProSST (Li et al., 2024). ProSST primarily improves upon SaProt (Su et al., 2023) by incorporating a larger vocabulary of structural tokens and employing disentangled attention mechanisms. Instead of relying on the 3Di alphabet optimized for protein structure search with Foldseek (van Kempen et al., 2022), Li et al. (2024) pre-train a new autoencoder to denoise corrupted protein backbones and cluster the resulting latent space using the $K$-means algorithm (Lloyd, 1982). Notably, optimal performance for fitness prediction is achieved with $K=2048$ tokens, compared to just 20 in the 3Di vocabulary used by SaProt. We adopt this model in our experiments. Additionally, disentangled attention in ProSST enhances information propagation between sequence and structure within its Transformer blocks, further improving prediction performance. The model has 110M parameters in total.
ProSST, similarly to ESM2 and SaProt, is pre-trained using masked language modeling applied to protein sequence tokens. To score mutations on the ProteinGym benchmark (Notin et al., 2024), ProSST also uses the log-odds ratio, but in a slightly different way compared to ESM2 and SaProt. Specifically, ProSST performs a single forward pass to predict log probabilities, which are then used to score all mutations. Formally, this approach modifies the log probability condition in Equation˜5, replacing $x_{\setminus i}$ with $x$.
ProSST + ProteinTTT.
Similarly to ESM2 and SaProt, we treat the Transformer encoder in ProSST as the backbone $f$, the masked language modeling head as the pre-training head $g$, and the log-odds ratio formula as the task-specific head $h$.
MSA Transformer.
Finally, we experiment with MSA Transformer for fitness prediction (Rao et al., 2021). Similar to ESM2 (Lin et al., 2023), MSA Transformer is pre-trained on large protein sequence datasets; however, it is trained on multiple sequence alignments (MSAs) rather than individual sequences.
Since MSA Transformer is also a protein language model, it can be used for fitness prediction in the same way as ESM2, as discussed above, by computing the log-odds ratio over the first sequence in the MSA in this case. We reproduce the results of MSA Transformer on the ProteinGym benchmark with two modifications: (1) we sample a weighted subset of 32 sequences from each MSA instead of 400, and (2) we use only one random seed instead of five for ensembling. These changes significantly reduce computational time while also slightly improving performance compared to the results reported in ProteinGym. This improvement may be explained by the fact that the performance of MSA Transformer saturates with increasing MSA input size (Figure 4 in Rao et al. (2021)).
MSA Transformer + ProteinTTT.
[…]
E.3 Protein function prediction
E.3.1 Datasets
@@ E.3.2 Metrics
The TPS substrate prediction problem is a 12-class multi-label classification task over possible TPS substrates. Therefore, we assess the quality of the predictions using standard multi-label classification metrics such as mean average precision (mAP) and area under the receiver operating characteristic curve (AUROC) averaged across individual classes. These metrics were used in the original EnzymeExplorer paper (Samusevich et al., 2025). We report the performance by averaging the metric values concatenated across all validation folds from the 5-fold cross-validation schema.
@@
E.3.3 Models
EnzymeExplorer.
EnzymeExplorer is a state-of-the-art method for the classification of terpene synthase (TPS) substrates (Samusevich et al., 2025). The model consists of two parallel tracks. Given a protein sequence, EnzymeExplorer first computes its ESM-1v embedding (Meier et al., 2021) and a vector of similarities to the functional domains of proteins from the training dataset, based on unsupervised domain segmentation of AlphaFold2-predicted structures (Jumper et al., 2021). The ESM-1v embedding and the similarity vector are then concatenated and processed by a separately trained random forest, which predicts TPS substrate class probabilities.
In our experiments, we use the “PLM only” version of the model, which leverages only ESM-1v embeddings. This version exhibits a minor performance decrease compared to the full model but exactly follows a Y-shaped architecture, allowing us to validate the effectiveness of ProteinTTT for predicting TPS substrates. We use the implementation of EnzymeExplorer available at the official GitHub page  [^10] https://github.com/pluskal-lab/EnzymeExplorer.
EnzymeExplorer + ProteinTTT.
When applying ProteinTTT to EnzymeExplorer, we treat the frozen ESM-1v model as a backbone $f$, its language modeling head as a self-supervised head $g$, and the random forest classifying TPS substrates as a downstream supervised head $h$.